OpenStudy (anonymous):

Robin bought a square sheet of paper of side 12 inches to decorate a pencil box he made in craft class. The pencil box has a height of 5 inches, a width of 3 inches, and a length of 4 inches. Approximately how many square inches of paper will be remaining if he pastes paper on to the bottom base and the four side surfaces of the pencil box without any overlap?

OpenStudy (anonymous):

please help

OpenStudy (jlastino):

I think i answered a question like this yesterday lol You should find the area to be covered on the pencil box (not including the top) lw+2wh+2lh Then find the area of the square sheet of paper A= s^2 Subtract the sum of the area on the pencil from the area of the sheet of paper

OpenStudy (anonymous):

so what's the answer

OpenStudy (jlastino):

Find the are that's going to cover the pencil case. 4(3)+2(3)(5)+2(4)(5) = 12+30+40 = 82 FInd the area of the sheet of paper = 12^2=144 Then subtract the area to be covered form the area of the paper 144-82 You will be left w/ 62 square inches